8th Nov Eve Shift Summary

KTeV Dart (da_ktev@ktev3)
Tue, 9 Nov 1999 00:10:58 -0600 (CST)

Shift Summary for eve shift, 8th Nov /99

On Shift: JW, RP

Hours of beam available at all intensities: 8hr
Hours of beam available at nominal intensity: 6.5hr
Hours of E799 data (special runs not included): 7hr
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:
beam intensity spotty - 1 turn .
pipeline hangup; restart problems; atev2 reboot.

Runs ended this shift:

Run 14929 - 10.6M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)


Accesses during shift:

Overview of shift:
E.g., calibrate CsI, E832 datataking

Description of Shift:

run spill time comment

14929 118 16.13 Beam gone for several spills
135 Lots of missed spills

17.10 Dave Northacker brought 2 new bottles of argon-ethane

172 17.25 ktsad.exe on atev2 not running. restart the process.

200 18.15 Phone MCR as per run managers requestto ask about SEM
being low. They say they are tuning the beam.

270 CSI 1654 no illumination. TRy to reset pipeline
20.00 Try to init pipeline but process hung. NFS mount errors
when try to reboot. JasonH notices atev2 swap space
error. BobT suggests reboot of Atev2.

20.15 Stop run. Reboot atev2.

14930 Update ADGN constants. Restart sdaq. Reboot pipeline
crates 12-16. Init pipeline with DBC. Start Run
KTevana errors. Channel 1654 is fine. All now ok.

20.43 Restart trd gas monitor.

13 20.50 MI power supply problems. estimate 15 mins.

21.00 Beam returns. RJT investigating L3 problem - finds
crate 32 has bad word order.

22.30 Missing beam for spills 86-91.

23.45 Onsplit tape mount - UPE008
23.55 Onsplit tape mount - UPH018

Follow up Notes on Previous Shift

Items to Follow up on Next Shift